OpenStudy (shubhamsrg):

Probability ques (easy yet worth trying) In an urn, there are several colored balls, with equal numbers of each color. If we were to add 14 balls of a new color (different from those in the urn) into the urn, the probability of drawing two balls, without replacement, of the same color stays constant. How many balls are there in the urn initially?

OpenStudy (sirm3d):

i'd say 14n, where n is the number of colors in the urn originally.

OpenStudy (shubhamsrg):

Actually there is a proper integer answer to this, and nopes,its not of the form 14n

OpenStudy (perl):

let k = # distinct colors n = # of balls per color , which is the same for each color. so there are a total of k*n balls , then you add 14 you get k*n + 14 balls , . now probability of taking two balls with replacement same color

OpenStudy (anonymous):

is it\[\frac{2}{n+14}\] for same two balls \[\frac{2k}{n}\] for color

OpenStudy (perl):

the probability before the 14 balls is put in is k* [n(n-1)] / ( k*n * (kn-1) ) the probability after the 14 balls for picking the same ball is (k+1)(n+1)(n) / ( nk + 14 )

OpenStudy (perl):

how about we do an example 4 red balls, 4 blue balls , 4 green balls. mix them up in an urn the probability of picking two balls of same color (w/out replacement) P( RR or BB or GG) = 3* 4/12 *3/11

OpenStudy (perl):

now add 14 yellow balls

OpenStudy (perl):

i get k*n(n-1) / [ nk * (nk-1) ] = [ k * n (n-1) + 182] / [ (nk + 14) ( nk + 13) ]

terenzreignz (terenzreignz):

By the way, there are 91 balls, methinks.

terenzreignz (terenzreignz):

And there are 7 balls per colour.

OpenStudy (perl):

terence, before or after the 14 are put in ?

OpenStudy (perl):

ok initially

OpenStudy (perl):

terence, your solution works in my equation :)

terenzreignz (terenzreignz):

Yay :) Also you bothered to spell "Terence" right XD

OpenStudy (perl):

k*n(n-1)/[ nk * (nk-1) ] = [ k *n (n-1) + 182] / [ (nk + 14) ( nk + 13) ]

OpenStudy (perl):

may i ask how you got your answer?

terenzreignz (terenzreignz):

Nothing too fancy, too much trial and error... at least I know the answer, now to find an efficient general way to solve this :>

OpenStudy (perl):

i dont see an easy way to solve my equation, i tried cross multiplying

OpenStudy (perl):

i dont consider this an easy question, the guy who posted this is torturing us

OpenStudy (anonymous):

yes i agree this is not an easy question @shubhamsrg wheres this from

terenzreignz (terenzreignz):

Tentatively, I think that Given an urn with n balls and k balls per colour, if adding r balls of a new colour would make it so that the probability of drawing two balls of the same colour remains constant, then there must have been rC2 balls initially, where C is the combination operator. So n = rC2

OpenStudy (shubhamsrg):

Well your answers are right. And seriously its not that hard. Its from brilliant.org level 4 ques. I was able to do it there somehow. I wished to share it here.

OpenStudy (shubhamsrg):

I should post my solution now ?

OpenStudy (perl):

please post

OpenStudy (shubhamsrg):

Let there be n balls, q sets where 1 set denotes 1 color, and k balls in each set, i.e. kq = n Now, probab of drawing 2 balls of same color initially can be seen as for 1st ball , we have n choices -> n/n = 1 for 2nd ball , we have only k-1 choices , -> (k-1)/(n-1) Hence probab for drawing 2 balls of same color initially = 1* (k-1)/n = (k-1)/(n-1) After 14 balls have been put in, lets consider 2 diff cases, 1) when 2 balls are drawn from original set of n balls 2) when 2 balls are drawn from these 14 balls probab of 1st case will be (n/(n+14)) * ((k-1)/(n+13)) probab of 2nd case will be (14/(n+14))* (13/(n+13)) Net probab = (n/(n+14)) * ((k-1)/(n+13)) + (14/(n+14))* (13/(n+13)) = (n(k-1) + 182)/((n+14)(n+13)) On equating, (k-1)/(n-1) = (n(k-1) + 182)/((n+14)(n+13)) (k-1)(n+13)(n+14) = (n(k-1) + 182)(n-1) (k-1)(n^2 + 27n + 182) = (n^2 -n)(k-1) + 182(n-1) kn^2 + 27kn + 182k -n^2 -27n -182 = n^2 k -n^2 -nk +n + 182n - 182 =>28kn + 182k -210n = 0 => 2nk + 13k - 15n = 0 => n(2k-15) = -13k n = (13k)/(15-2k) Hit and trial for (n,k) gives (1,1) <-- can not be solution since 2 balls are drawn (13,5) (26,6) (91,7) Of these only (91,7) is a solution since n/k for remaining 2 isnt integer (recall n/k = q)

OpenStudy (anonymous):

oh teri..... :O